OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-2, 0) and a directrix at x = 2.

OpenStudy (antonio_xx2):

x^2 = 4ay = 4(2)y = 8y. Rearranging gives: y = 1/8*x^2

OpenStudy (anonymous):

ahhhhhh i see it

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

but wouldnt there be a negative since its -2

OpenStudy (antonio_xx2):

idk u tell me;)

OpenStudy (anonymous):

YES THERE IS

OpenStudy (anonymous):

THANKS BRUHHH <3

OpenStudy (antonio_xx2):

sure

OpenStudy (amistre64):

isnt this of the form: x = y^2 ?

OpenStudy (anonymous):

so y and x would be switched

OpenStudy (anonymous):

x = 1/8*y^2

OpenStudy (amistre64):

theyd have to be in order for a directrix to be a line x=2 

OpenStudy (anonymous):

ahhhhhh